OpenStudy (anonymous):

Solve for W: 2w^3-w^2+6w-3=0

OpenStudy (anonymous):

factor by grouping and you'll get ^_^:\[= w^2(2w-1) + 3(2w-1) = (2w-1)(w^2+3)\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

you can also use horner's method of factoring polynomial

OpenStudy (amistre64):

not to be a stickler in the mud; but there is no "W" in the problem :)

OpenStudy (anonymous):

lol, he meant w :)

OpenStudy (amistre64):

meant, schmeant.... lol

OpenStudy (anonymous):

lol, alright ^_^

OpenStudy (anonymous):

Not to be a stickler, but the phrase is, "a stick in the mud"

OpenStudy (anonymous):

\[\log_{17} 17^{y}\]

OpenStudy (anonymous):

Remember that \[k = log_{b}x \iff b^k = x\]

OpenStudy (anonymous):

so let \[ k = log_{17} 17^y \rightarrow 17^? = 17^y\]

OpenStudy (amistre64):

canadian eh....

OpenStudy (anonymous):

\[\log_{17} 17^{y}\]

OpenStudy (anonymous):

... are you repeating the question for any particular reason?

OpenStudy (anonymous):

Nah, i got it. typo. Thanks anyway